If Mark is unable to see objects clearly when they are placed beyond 0.5m away, (a) is he near-sighted or far-sighted? (b) What should be the focal length of a pair of contact lenses prescribed to correct this problem? (c) If instead he wears eye-glasses, and they are to sit 2.0cm in front of his eyes, what prescription should they be?
Answer #1

He is near sighted.

we know that focal length equation is,

1/f=1/u + 1/v where,

u=infinity so 1/f=1/v=-1/50cm

i.e. f=-50cm